The cupcake café makes 4 1/2 times as much revenue on doughnuts as muffins. If total sales were $44,000 for May, what dollar amount of each was sold?
Answer by TimothyLamb(4379)

x = donut sales
y = muffin sales
---
x = 4.5y
x + y = 44000
---
put the system of linear equations into standard form
---
x - 4.5y = 0
x + y = 44000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = donut sales = 36000
y = muffin sales = 8000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php